OpenStudy (anonymous):

Integrate: (2e^x+2/x^(1/2)+3)dx

OpenStudy (anonymous):

go to wolframalpha.com and type this in

OpenStudy (anonymous):

\[\int\limits_{}^{}(3e ^{x}+2/\sqrt{x}+3)dx=\int\limits_{}^{}e ^{x}dx+\int\limits_{}^{}2/\sqrt{x}dx+\int\limits_{}^{}3dx\] this gives you \[3e ^{x}+4\sqrt{x}+3x+c\] where c is the integration constant.